# Efficient and Scalable Density Functional Theory Hamiltonian Prediction through Adaptive Sparsity

### v1

@@ Abstract
Hamiltonian matrix prediction is pivotal in computational chemistry, serving as the foundation for determining a wide range of molecular properties. While SE(3) equivariant graph neural networks have achieved remarkable success in this domain, their substantial computational cost—driven by high-order tensor product (TP) operations—restricts their scalability to large molecular systems with extensive basis sets. To address this challenge, we introduce SPHNet, an efficient and scalable equivariant network, that incorporates adaptive SParsity into Hamiltonian prediction. SPHNet employs two innovative sparse gates to selectively constrain non-critical interaction combinations, significantly reducing tensor product computations while maintaining accuracy. To optimize the sparse representation, we develop a Three-phase Sparsity Scheduler, ensuring stable convergence and achieving high performance at sparsity rates of up to 70%. Extensive evaluations on QH9 and PubchemQH datasets demonstrate that SPHNet achieves state-of-the-art accuracy while providing up to a 7x speedup over existing models. Beyond Hamiltonian prediction, the proposed sparsification techniques also hold significant potential for improving the efficiency and scalability of other SE(3) equivariant networks, further broadening their applicability and impact.
@@ 1 Introduction
The Kohn-Sham Hamiltonian matrix is a fundamental quantity in computational chemistry, as it enables the prediction of all molecular properties obtainable through traditional Density Functional Theory (DFT) calculations (Hohenberg & Kohn, 1964; te Vrugt et al., 2020). Accurately predicting the Hamiltonian matrix allows for the rapid determination of system energy, HOMO-LUMO gaps, electron density, and other important properties (Fang et al., 2022; Zang et al., 2023; Batzner et al., 2022; Batatia et al., 2022; Wang et al., 2022; Li et al., 2024; Chen et al., 2023). Consequently, recent efforts have focused on using machine learning techniques to predict the Hamiltonian matrix (Schütt et al., 2017; Unke et al., 2021; Yu et al., 2023; Gong et al., 2023; Zhou et al., 2022; Zhong et al., 2023). Among these, methods based on SE(3) equivariant graph neural networks (Unke et al., 2021; Yu et al., 2023; Gong et al., 2023; Yu et al., 2023; Li et al., 2025) have achieved superior accuracy due to their consistency with the equivariance of the Hamiltonian matrix. These methods typically project molecular features into spherical harmonics space and use tensor product operations to interact features of different orders, which are crucial for maintaining equivariance.
@@ 2 Related Works
Hamiltonian Matrix Prediction. Hamiltonian prediction has advanced with neural networks in recent years. SchNOrb(Schütt et al., 2019) extended SchNet (Schütt et al., 2017) for high-accuracy molecular orbital predictions, while PhiSNet (Unke et al., 2021) successfully introduced SE(3) networks, significantly improving accuracy while facing inefficiency due to tensor product operations. QHNet (Yu et al., 2023) improved efficiency by reducing the number of tensor products and (Li et al., 2025) further extends the scalability and applicability of Hamiltonian matrix prediction on large molecular systems by introducing a novel loss function. At the same time, methods such as DeepH (Li et al., 2022) are focused on the Hamiltonian prediction of the periodic system.
@@ 3 Preliminary
where $\mathbf{C}$ representing the coefficients of molecular orbitals, $\mathbf{H}$ signifies the Hamiltonian matrix, $\mathbf{S}$ refers to the overlap matrix, and $\boldsymbol{\epsilon}$ is the diagonal matrix encapsulating orbital energies. Note that $\mathbf{C}\in\mathbb{R}^{n\times n}$ where $n$ denotes the number of basis functions and $\mathbf{H},\mathbf{S},\boldsymbol{\epsilon}\in\mathbb{R}^{n_{o}\times n_{o}}$ where $n_{o}$ denotes the number of atomic orbitals used in DFT calculations. Primarily, an iterative self-consistent field (SCF) methodology (Payne et al., 1992; Cances & Le Bris, 2000; Wang & Baerends, 2022; Kudin et al., 2002) is utilized to solve the Kohn-Sham (KS) equations and to obtain the Hamiltonian matrix, requiring a temporal complexity of $\mathcal{O}(n^{3})$.
@@ 4.1 Three-phase Sparsity Scheduler
Specifically, in the first phase, $\mathbf{W}$ is initialized as an all-ones vector, and $\text{RANDOM}(\cdot)$ denotes a random function that selects elements from the set independently of $\mathbf{W}$ with a given probability $1-k$. This ensures that all parameters are unbiasedly selected and updated. In the second phase, $\text{TOP}(\cdot)$ is used to select the elements with the highest weights from the set with a given probability $1-k$, ensuring that the optimal combinations are chosen globally. In the third phase, ${\mathbf{W}^{p_{2}}}^{\prime}$ represents a fixed vector without gradients, which is frozen after the final backward update of the second phase. From this point onward, the selected combinations remain unchanged. This design prioritizes efficiency, as static connections typically result in faster computation.
@@ 4.3 Sparse Pair Gate
$\mathbf{I}_{ij}=(\mathbf{x}^{0}_{i}\|\mathbf{x}^{0}_{j}\|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle^{0}),$ (6)
@@
where $F_{r}(\cdot)$ and $F_{s}(\cdot)$ are linear layers, $\mathrm{RBF}(\cdot)$ is the radial basis function.
@@ 5 Result
We conducted a series of experiments to compare the overall performance of SPHNet with the previous state-of-the-art model, including SchNOrb(Schütt et al., 2019), PhiSNet(Unke et al., 2021), QHNet(Yu et al., 2023) and WANet(Li et al., 2025), and the results listed in the table are sourced from the their papers. Note that the results for PhiSNet and SchNOrb can only be conducted on the MD17 dataset, as these models are specifically designed for trajectory datasets in conformational space. Additionally, since WANet has not been open-sourced, we only report results on a subset of datasets and metrics based on the information provided in their paper. Detailed experimental settings are described in Appendix A. Additional ablation studies on the Three-phase Sparsity Scheduler and Interaction Block, as well as further analysis of the experiments presented in this section, are provided in Appendix B.
Datasets. Three datasets were used in the experiments: The MD17 dataset includes Hamiltonian matrices for the trajectory data of four molecules: water, ethanol, malondialdehyde, and uracil, containing 4,900, 30,000, 26,978, and 30,000 structures, respectively. The QH9 dataset (Yu et al., 2024), consisting of Hamiltonian matrices for 134k small molecules with no more than 20 atoms, and the PubChemQH dataset (Li et al., 2025), containing Hamiltonian matrices for 50k molecules with atom counts ranging from 40 to 100. The QH9 and PubChemQH use the B3LYP exchange-correlation function, and the MD17 uses the PBE exchange-correlation function. The GTO orbital basis for MD17 and QH9 is Def2-SVP while PubChemQH uses the Def2-TZVP GTO orbital basis. Compared to the MD17 and QH9 datasets, the maximum number of orbitals in the Hamiltonian of PubChemQH dataset increases by about 15 times, resulting in a 200x difference in the number of elements in the Hamiltonian matrix.
@@ 5.1 Performance on QH9 Dataset
We first evaluated the performance of SPHNet on the QH9 dataset. Noted that the QH9 dataset has two subsets and four different splits, including the stable-iid, stable-ood, dynamic-geo, and dynamic-mol. We trained SPHNet on four different split sets and compared the results with the baseline models QHNet and WANet. As shown in Table 1, for two stable split sets, SPHNet attained a speedup of 3.7x to 4x over QHNet while improving the accuracy of the Hamiltonian Mean Absolute Error (MAE) and the occupied energy MAE. Furthermore, SPHNet significantly reduced GPU memory usage, requiring only 30% memory usage of the baseline model. For two dynamic split sets, SPHNet maintained a speedup of 3.3x to 3.5x, simultaneously enhancing prediction accuracy and decreasing GPU memory usage by over 70%. These results underscore SPHNet’s efficiency, achieving substantial speedups and resource savings without sacrificing precision, indicating that SPHNet can effectively learn molecular features and the Hamiltonian matrix within the small-scale dataset, establishing a solid foundation for calculations in larger-scale systems.
@@ 5.2 Performance on PubChemQH Dataset
We observed that SPHNet achieves a higher speedup on the PubChemQH dataset. As described in the dataset section, the number of orbitals in the Hamiltonians of PubChemQH is 15 times greater than in QH9. Additionally, since PubChemQH was computed using the Def2-TZVP orbital basis, the network requires a maximum output angular momentum order of $L_{\text{max}}=6$, whereas datasets computed with the Def2-SVP basis only require $L_{\text{max}}=4$. These make the number and time consumption of tensor products much higher than that in small-scale datasets. Thus, we can reasonably infer that the higher speedup achieved on the PubChemQH dataset is due to the markedly greater computational complexity of its molecules compared to those in the MD17 and QH9 datasets. This increased complexity introduces higher information redundancy, making it more amenable to adaptive sparsification. Our sparsity ablation experiments across different datasets support this inference. As molecular size and maximum angular momentum order increase, the rising sparsity rates of sparse pair gates and sparse tensor product gates generally result in smaller accuracy losses. Therefore, higher coefficient rates can be applied to SPHNet, further accelerating its performance on large-scale datasets. For more details on this ablation study, please refer to Section 5.4.
@@ 5.4 Ablation Study on Sparsity Rate
The key idea of SPHNet is the implementation of adaptive sparsity to reduce computational demands in SE(3) equivariant networks. The most critical hyperparameter in this adaptive sparse system is the sparsity rate. To determine the optimal sparsity rate and assess its impact across various molecular systems, we conducted a series of experiments. We varied the sparsity rate from 0% to 90% in 10% intervals and trained SPHNet on three datasets representing different molecular scales (Here we chose Ethanol to represent MD17 dataset). As illustrated in Fig.3, the predicted Hamiltonian MAE across all three datasets remained relatively stable at lower sparsity rates but increased significantly when the rate reached a specific threshold. This finding suggests that an appropriate range of sparsity has minimal impact on model accuracy while simultaneously enhancing computational speed.
Figure 3: The effect of different sparsity rates on the model performance on the datasets with different scale of molecules.
@@ 5.5 Analysis of Optimal Selected Set
Figure 4: (A) The distribution of distance between selected node pairs and the selected proportions of different pair distances in the pair gate. (B) The proportions of each output order being selected in the first Non-Diagonal block’s tensor product gate.
@@ 6 Conclusion and Future Work
This study opens several promising avenues for future work. First, the adaptive sparsity technique shows considerable potential for generalization to a wider range of tasks. Specifically, the sparse tensor product gates can be readily extended to any SE(3)-equivariant network architecture based on Clebsch-Gordan tensor products, while the sparse pair gates could be adapted to more types of pairwise interactions, such as those found in MACE (Batatia et al., 2022) for multi-body interactions. Second, integrating advanced loss functions, such as those introduced in (Li et al., 2025), could further improve the accuracy of downstream properties of the Hamiltonian matrix, particularly in enhancing the applicability for large molecular systems. These directions will guide future efforts with further experimental validation, expanding the impact of adaptive sparsification and optimizing the performance of SE(3)-equivariant networks.

### v2

@@ Abstract
Hamiltonian matrix prediction is pivotal in computational chemistry, serving as the foundation for determining a wide range of molecular properties. While SE(3) equivariant graph neural networks have achieved remarkable success in this domain, their substantial computational cost—driven by high-order tensor product (TP) operations—restricts their scalability to large molecular systems with extensive basis sets. To address this challenge, we introduce SPHNet, an efficient and scalable equivariant network, that incorporates adaptive SParsity into Hamiltonian prediction. SPHNet employs two innovative sparse gates to selectively constrain non-critical interaction combinations, significantly reducing tensor product computations while maintaining accuracy. To optimize the sparse representation, we develop a Three-phase Sparsity Scheduler, ensuring stable convergence and achieving high performance at sparsity rates of up to 70%. Extensive evaluations on QH9 and PubchemQH datasets demonstrate that SPHNet achieves state-of-the-art accuracy while providing up to a 7x speedup over existing models. Beyond Hamiltonian prediction, the proposed sparsification techniques also hold significant potential for improving the efficiency and scalability of other SE(3) equivariant networks, further broadening their applicability and impact. Our code can be found at https://github.com/microsoft/SPHNet.
@@ 1 Introduction
The Kohn-Sham Hamiltonian matrix is a fundamental quantity in computational chemistry, as it enables the prediction of all molecular properties obtainable through traditional Density Functional Theory (DFT) calculations (Hohenberg & Kohn, 1964; te Vrugt et al., 2020). Accurately predicting the Hamiltonian matrix allows for the rapid determination of system energy, HOMO-LUMO gaps, electron density, and other important properties (Fang et al., 2022; Zang et al., 2023; Batzner et al., 2022; Batatia et al., 2022; Wang et al., 2022; Li et al., 2024; Chen et al., 2024). Consequently, recent efforts have focused on using machine learning techniques to predict the Hamiltonian matrix (Schütt et al., 2017; Unke et al., 2021; Yu et al., 2023; Gong et al., 2023; Zhou et al., 2022; Zhong et al., 2023). Among these, methods based on SE(3) equivariant graph neural networks (Unke et al., 2021; Yu et al., 2023; Gong et al., 2023; Yu et al., 2023; Li et al., 2025b) have achieved superior accuracy due to their consistency with the equivariance of the Hamiltonian matrix. These methods typically project molecular features into spherical harmonics space and use tensor product operations to interact features of different orders, which are crucial for maintaining equivariance.
@@ 2 Related Works
Hamiltonian Matrix Prediction. Hamiltonian prediction has advanced with neural networks in recent years. SchNOrb (Schütt et al., 2019) extended SchNet (Schütt et al., 2017) for high-accuracy molecular orbital predictions, while PhiSNet (Unke et al., 2021) successfully introduced SE(3) networks, significantly improving accuracy while facing inefficiency due to tensor product operations. QHNet (Yu et al., 2023) improved efficiency by reducing the number of tensor products and (Li et al., 2025b) further extends the scalability and applicability of Hamiltonian matrix prediction on large molecular systems by introducing a novel loss function. At the same time, methods such as DeepH (Li et al., 2022) are focused on the Hamiltonian prediction of the periodic system.
@@ 3 Preliminary
where $\mathbf{C}$ representing the coefficients of molecular orbitals, $\mathbf{H}$ signifies the Hamiltonian matrix, $\mathbf{S}$ refers to the overlap matrix, and $\boldsymbol{\epsilon}$ is the diagonal matrix encapsulating orbital energies. Note that $\mathbf{C}\in\mathbb{R}^{n\times n_{0}}$, $\mathbf{H},\mathbf{S}\in\mathbb{R}^{n\times n}$, and $\boldsymbol{\epsilon}\in\mathbb{R}^{n_{o}\times n_{o}}$, where $n$ denotes the number of basis functions and $n_{o}$ denotes the number of atomic orbitals used in DFT calculations. Primarily, an iterative self-consistent field (SCF) methodology (Payne et al., 1992; Cances & Le Bris, 2000; Wang & Baerends, 2022; Kudin et al., 2002) is utilized to solve the Kohn-Sham (KS) equations and to obtain the Hamiltonian matrix, requiring a temporal complexity of $\mathcal{O}(n^{3})$.
@@ 4.1 Three-phase Sparsity Scheduler
Specifically, in the first phase, $\mathbf{W}$ is initialized as an all-ones vector, and $\text{RANDOM}(\cdot)$ denotes a random function that selects elements from the set independently of $\mathbf{W}$ with a given probability $1-k$. This ensures that all parameters are unbiasedly selected and updated. In the second phase, $\text{TOP}(\cdot)$ is used to select the elements whose weights are within the top $1-k$ percent of all elements, ensuring that the optimal combinations are chosen globally. In the third phase, ${\mathbf{W}^{p_{2}}}^{\prime}$ represents a fixed vector without gradients, which is frozen after the final backward update of the second phase. From this point onward, the selected combinations remain unchanged. This design prioritizes efficiency, as static connections typically result in faster computation.
@@ 4.3 Sparse Pair Gate
$\mathbf{I}_{ij}=(\mathbf{x}^{0}_{i}\|\mathbf{x}^{0}_{j}\|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle^{1:}),$ (6)
@@
where $F_{r}(\cdot)$ and $F_{s}(\cdot)$ are linear layers, $\mathrm{RBF}(\cdot)$ is the radial basis function. The extra computational cost introduced by above sparse gates and the scheduler is minimal and has negligible impact on the overall time, which is discussed in detailed in Appendix D.7.
@@ 5 Result
We conducted a series of experiments to compare the overall performance of SPHNet with the previous state-of-the-art model, including SchNOrb (Schütt et al., 2019), PhiSNet (Unke et al., 2021), QHNet (Yu et al., 2023), and WANet (Li et al., 2025b), and the results listed in the table are sourced from their papers. Note that the results for PhiSNet and SchNOrb can only be conducted on the MD17 dataset, as these models are specifically designed for trajectory datasets in conformational space. Additionally, since WANet has not been open-sourced, we only report results on a subset of datasets and metrics based on the information provided in their paper. Detailed experimental settings are described in Appendix A.
Datasets. Three datasets were used in the experiments: The MD17 dataset includes Hamiltonian matrices for the trajectory data of four molecules: water, ethanol, malondialdehyde, and uracil, containing 4,900, 30,000, 26,978, and 30,000 structures, respectively. The QH9 dataset (Yu et al., 2024), consisting of Hamiltonian matrices for 134k small molecules with no more than 20 atoms, and the PubChemQH dataset (Li et al., 2025b), containing Hamiltonian matrices for 50k molecules with atom counts ranging from 40 to 100. The QH9 and PubChemQH use the B3LYP exchange-correlation function, and the MD17 uses the PBE exchange-correlation function. The orbital basis Def2-SVP is used for MD17 and QH9, while the Def2-TZVP is used for PubChemQH. Compared to the MD17 and QH9 datasets, the maximum number of orbitals in the Hamiltonian of the PubChemQH dataset increases by about 15 times, resulting in a 200x difference in the number of elements in the Hamiltonian matrix.
@@ 5.1 Performance on QH9 Dataset
We first evaluated the performance of SPHNet on the QH9 dataset. Note that the QH9 dataset has two subsets and four different splits, including the stable-iid, stable-ood, dynamic-geo, and dynamic-mol. We trained SPHNet on four different split sets and compared the results with the baseline models QHNet and WANet. As shown in Table 1, for two stable split sets, SPHNet attained a speedup of 3.7x to 4x over QHNet while improving the accuracy of the Hamiltonian Mean Absolute Error (MAE) and the occupied energy MAE. Furthermore, SPHNet significantly reduced GPU memory usage, requiring only 30% memory usage of the baseline model. For two dynamic split sets, SPHNet maintained a speedup of 3.3x to 3.5x, simultaneously enhancing prediction accuracy and decreasing GPU memory usage by over 70%. These results underscore SPHNet’s efficiency, achieving substantial speedups and resource savings without sacrificing precision, indicating that SPHNet can effectively learn molecular features and the Hamiltonian matrix within the small-scale dataset, establishing a solid foundation for calculations in larger-scale systems.
@@ 5.2 Performance on PubChemQH Dataset
We observed that SPHNet achieves a higher speedup on the PubChemQH dataset. As described in the dataset section, the number of orbitals in the Hamiltonians of PubChemQH is 15 times greater than in QH9. Additionally, since PubChemQH was computed using the Def2-TZVP orbital basis, the network requires a maximum output angular momentum order of $L_{\text{max}}=6$, whereas datasets computed with the Def2-SVP basis only require $L_{\text{max}}=4$. These make the number and time consumption of tensor products much higher than that in small-scale datasets. Thus, we can reasonably infer that the higher speedup achieved on the PubChemQH dataset is due to the markedly greater computational complexity of its molecules compared to those in the MD17 and QH9 datasets. This increased complexity introduces higher information redundancy, making it more amenable to adaptive sparsification. Our sparsity ablation experiments across different datasets support this inference. As molecular size and maximum angular momentum order increase, the rising sparsity rates of sparse pair gates and sparse tensor product gates generally result in smaller accuracy losses. Therefore, higher sparsity rates can be applied to SPHNet, further accelerating its performance on large-scale datasets. For more details on this ablation study, please refer to Section 5.4.
@@ 5.4 Ablation Study on Sparsity Rate
The key idea of SPHNet is the implementation of adaptive sparsity to reduce computational demands in SE(3) equivariant networks. The most critical hyperparameter in this adaptive sparse system is the sparsity rate. To determine the optimal sparsity rate and assess its impact across various molecular systems, we conducted a series of experiments. We varied the sparsity rate from 0% to 90% in 10% intervals and trained SPHNet on three datasets representing different molecular scales (Here we chose Ethanol to represent MD17 dataset.). As illustrated in Fig.3, the predicted Hamiltonian MAE across all three datasets remained relatively stable at lower sparsity rates but increased significantly when the rate reached a specific threshold. This finding suggests that an appropriate range of sparsity has minimal impact on model accuracy while simultaneously enhancing computational speed.
Figure 3: The effect of different sparsity rates on the model performance on the datasets with different scale of molecules. See the detailed computational cost scaling in Appendix B.2 of PubChemQH Dataset.
@@ 5.5 Analysis of Optimal Selected Set
Figure 4: (A) The distribution of distance between selected node pairs and the selected proportions of different pair distances in the pair gate. (B) The proportions of each output order being selected in the first Non-Diagonal block’s tensor product gate. See the full cases of selected pairs in Appendix B.4.
@@ 5.6 Scaling with Increasing Size of Hamiltonian
5.7 Additional Ablation study on SPHNet
To evaluate the effectiveness of the Three-phase Sparsity scheduler and the Sparse Gate on the model performance, we carried out a series of ablation studies. The results showed that these strategies have significant improvements in model efficiency and accuracy. The details are as follows.
Three-phase Sparsity Scheduler. We evaluated the effect of the three-phase sparsity scheduler on model performance by training SPHNet under various ablation settings: excluding the random stage, the adaptive stage, or the fixed stage. Each configuration was trained five times to assess performance stability and accuracy. As shown in Table 4, models trained without the random stage often converged to local optima, resulting in suboptimal performance. Models trained without the adaptive stage exhibited greater variance in outcomes and significantly lower accuracy compared to SPHNet with the full scheduler. Although omitting the fixed stage did not substantially affect accuracy, it reduced the speedup ratio to 5.45 due to the additional computational overhead from the sparse gate and the non-static graph implementation of tensor products following the Sparse Pair Gate, in contrast to the optimized implementations in the e3nn library. These findings highlight the importance of the three-phase sparsity scheduler in enhancing model performance and efficiency.
Table 4: Ablation Study on Three-phase Sparsity Scheduler.
Random Stage
Adaptive Stage
Fixed Stage
$H$ [$10^{-6}E_{h}$] $\downarrow$ $97.31$ $112.68$ $122.79$ $97.11$
±$\pm 0.52$$\pm 10.75$$\pm 19.02$ $1.33$
Memory [GB/Sample] $\downarrow$ 5.62 5.62 5.62 5.62
Speed [Sample/Sec] $\uparrow$ 3.12 3.12 3.12 3.12
Speedup Ratio $\uparrow$ 7.09× 7.09× 7.09× 5.45×
Sparse Gates. To further evaluate the impact of the Sparse Pair Gate and the Sparse Tensor Product Gate on overall performance, we trained SPHNet under three configurations: without the Sparse Pair Gate, without the Sparse Tensor Product Gate, and without both gates. As shown in Table 5, the Sparse Pair Gate achieved a 78% speedup and a 30% reduction in GPU memory usage, with minimal impact on model accuracy. The Sparse Tensor Product Gate, which removes 70% of the combinations in the tensor product, yielded a 160% speedup with an acceptable loss in accuracy. These results underscore the contribution of the Sparse Gate to improving model efficiency. Users may refer to Section 5.4 for guidance on choosing a sparsity ratio that balances speed and accuracy based on system size.
Table 5: Ablation Study on Sparse Gates.
Sparse Pair Gate
Sparse TP Gate
$H$ [$10^{-6}E_{h}$] $\downarrow$ 97.31 94.31 87.70 86.35
Memory [GB/Sample] $\downarrow$ 5.62 8.04 6.98 10.91
Speed [Sample/Sec] $\uparrow$ 3.12 1.75 1.20 0.76
Speedup Ratio $\uparrow$ 7.09× 3.98× 2.73× 1.73×
Applying Sparse Gate to QHNet model. To further assess the effectiveness of the Sparse Gates, we conducted additional experiments by integrating them into the QHNet model. Specifically, the Sparse Pair Gate was applied to the second non-diagonal pair block, while the Sparse Tensor Product Gate was applied to the node-wise interaction blocks as well as both diagonal and non-diagonal pair blocks. As shown in the table below, the sparse gates significantly improved the training speed of the QHNet model and reduced computational resource usage. These results demonstrate the generalizability of sparse gating mechanisms across different SE(3)-equivariant networks.
Table 6: The effect of sparse gates on QHNet model.
Sparse Pair Gate
Sparse TP Gate
$H$ [$10^{-6}E_{h}$] $\downarrow$ 123.74 128.16 126.27 128.89
Memory [GB/Sample] $\downarrow$ 22.50 12.68 10.07 8.46
Speed [Sample/Sec] $\uparrow$ 0.44 0.90 0.73 1.45
Speedup Ratio $\uparrow$ 1.00× 2.04× 1.66× 3.30×
@@ 6 Conclusion and Future Work
This study opens several promising avenues for future work. First, the adaptive sparsity technique shows considerable potential for generalization to a wider range of tasks. Specifically, the sparse tensor product gates can be readily extended to any SE(3)-equivariant network architecture based on Clebsch-Gordan tensor products, while the sparse pair gates could be adapted to more types of pairwise interactions, such as those found in MACE (Batatia et al., 2022) for multi-body interactions. Second, integrating advanced loss functions, such as those introduced in (Li et al., 2025b), could further improve the accuracy of downstream properties of the Hamiltonian matrix, particularly in enhancing the applicability for large molecular systems. These directions will guide future efforts with further experimental validation, expanding the impact of adaptive sparsification and optimizing the performance of SE(3)-equivariant networks.
Impact Statement
This paper presents work whose goal is to speed up the Hamiltonian matrix prediction process and advance the SE(3) equivariant network. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.